# Associative memory and dead neurons

### v1

@@ 4 Associative memories without dead neurons
To avoid these difficulties we will slightly modify the dynamics of associative memory (1) to minimally alter steady states and get a better energy function. The resulting system that we propose reads:
@@
If one assumes that it is possible to select $\boldsymbol{R}(\boldsymbol{u})\neq 0$ for all $\boldsymbol{u}$, the condition that we will verify later, steady states of the newly introduced system (9) follows from the old one with affine transformation $\boldsymbol{y}=\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}$. This means all architectures possible with the old model (1) are also possible with the new one (9).
@@
$\begin{split}&E(\boldsymbol{u};\alpha,\beta,\gamma,\boldsymbol{S})=\alpha E_{1}(\boldsymbol{u})+\beta E_{2}(\boldsymbol{u})+\gamma E_{3}(\boldsymbol{u};\boldsymbol{S}),\alpha,\beta,\gamma\in\mathbb{R},\\ &E_{1}(\boldsymbol{u})=\left(\boldsymbol{u}^{\top}\boldsymbol{W}\boldsymbol{g}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)-L\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)-\frac{1}{2}\boldsymbol{g}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)^{\top}\boldsymbol{W}\boldsymbol{g}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)\right),\\ &E_{2}(\boldsymbol{u})=\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{W}\boldsymbol{u}-L\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right),\,E_{3}\left(\boldsymbol{u};\boldsymbol{S}\right)=\frac{1}{2}\left(\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b})\right)^{\top}\boldsymbol{S}\left(\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b})\right)\end{split}$ (10)
@@
    $E(\boldsymbol{u};\alpha,\beta,\gamma,\boldsymbol{S})$ is a Lyapunov function for dynamical system (9) if one can find $\boldsymbol{Q}\geq 0$, $\boldsymbol{R}(\boldsymbol{u})$, $\boldsymbol{S}=\boldsymbol{S}^{\top}$ such that
    $\boldsymbol{R}^{\top}\boldsymbol{F}+\boldsymbol{F}^{\top}\boldsymbol{R}\geq\boldsymbol{Q},\boldsymbol{F}(\boldsymbol{u})=\left(\boldsymbol{W}\boldsymbol{\Lambda}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)\left(\alpha\boldsymbol{W}-\gamma\boldsymbol{S}\right)+\beta\boldsymbol{W}+\gamma\boldsymbol{S}\right)$ (11)
    holds for all $\boldsymbol{u}$.
@@
Proof.
  1. 1.
    Using the identities above we find
    $\begin{split}&\frac{\partial E_{1}}{\partial\boldsymbol{u}}=\boldsymbol{W}\boldsymbol{\Lambda}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)\boldsymbol{W}\left(\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}-\boldsymbol{b})\right),\,\frac{\partial E_{2}}{\partial\boldsymbol{u}}=\boldsymbol{W}\left(\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}-\boldsymbol{b})\right),\\ &\frac{\partial E_{3}}{\partial\boldsymbol{u}}=\left(\boldsymbol{I}-\boldsymbol{W}\boldsymbol{\Lambda}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)\right)\boldsymbol{S}\left(\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}-\boldsymbol{b})\right).\\ \end{split}$
    Multiplying gradients on $\alpha,\beta,\gamma$ and adding them gives $\boldsymbol{F}(\boldsymbol{u})$. Since $\dot{\boldsymbol{u}}=\boldsymbol{R}(\boldsymbol{u})\boldsymbol{f}(\boldsymbol{u}(t))$ where $\boldsymbol{f}(\boldsymbol{u}(t))=\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}-\boldsymbol{b})$ from $\dot{E}=\dot{\boldsymbol{u}}^{\top}\frac{\partial E}{\partial\boldsymbol{u}}$ we obtain
  2. 2.
  3. 3.
  4. 4.
@@
In this case $E=E_{1}$ condition (11) becomes $\boldsymbol{R}^{\top}\boldsymbol{W}+\boldsymbol{W}\boldsymbol{R}\geq 0$. There are many strategies to select $\boldsymbol{R}$ and $\boldsymbol{W}$:
@@
Here we build a concrete example of hierarchical associative memory [15] with smooth Leaky ReLU [1] to avoid dealing with ODE having a non-smooth right-hand side. The activation function reads $\boldsymbol{g}(\boldsymbol{u})=\boldsymbol{u}\left(5+3\text{erf}\left(3\boldsymbol{u}\big{/}8\right)\right)\big{/}8$ and its derivative is $\boldsymbol{g}^{{}^{\prime}}(\boldsymbol{u})=\boldsymbol{g}(\boldsymbol{u})\big{/}\boldsymbol{u}+9\boldsymbol{u}\exp\left(-9\boldsymbol{u}^{2}\big{/}64\right)\big{/}\left(32\sqrt{\pi}\right)$, where all operations are pointwise and erf is error function. Weights and state vector are selected as in Example 1 but without requirement $\boldsymbol{W}_{ii+1}=\boldsymbol{W}_{i+1i}^{\top}$. After that we select $\boldsymbol{R}=\boldsymbol{I}-\boldsymbol{W}^{\top}\boldsymbol{\Lambda}$ and obtain the following dynamical system
@@
For this system Lyapunov function
$E_{2}=\frac{1}{2}\left(\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}(t)+\boldsymbol{b})-\boldsymbol{u}(t)\right)^{\top}\left(\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}(t)+\boldsymbol{b})-\boldsymbol{u}(t)\right)$
is non-increasing on trajectories. Matrix $\left(\boldsymbol{I}-\boldsymbol{W}^{\top}\odot\boldsymbol{g}^{{}^{\prime}}(\boldsymbol{W}\boldsymbol{u}(t)+\boldsymbol{b})\right)$ can in principle have non-trivial nullspace. To avoid this one observe that $\left\|\boldsymbol{W}\right\|_{2}\leq\max_{i}\left\|\boldsymbol{W}_{ii+1}\right\|_{2}+\max_{i}\left\|\boldsymbol{W}_{ii-1}\right\|_{2}$ (the bound is tight) and take $\boldsymbol{W}_{ii+1}=\widetilde{\boldsymbol{W}}_{ii\pm 1}\big{/}\left(2\left\|\widetilde{\boldsymbol{W}}_{ii\pm 1}\right\|_{2}\right)$, for some $\widetilde{\boldsymbol{W}}$ with the same block structure. Note, that in this construction $\boldsymbol{\Lambda}$ is not positive definite, and $\boldsymbol{W}$ is not symmetric.
All theoretical statements that we derived are valid only for modified dynamical system 9. However, it is easy to see that energy function $E_{3}$ from Proposition 5 can be adopted for the original temporal dynamics (1). The resulting energy function and dynamical system reads
@@
$E_{3}(\boldsymbol{u})=\frac{1}{2}\left(\boldsymbol{W}\boldsymbol{g}(\boldsymbol{y}(t))-\boldsymbol{y}(t)+\boldsymbol{b}\right)^{\top}\boldsymbol{S}\left(\boldsymbol{W}\boldsymbol{g}(\boldsymbol{y}(t))-\boldsymbol{y}(t)+\boldsymbol{b}\right).$ (14)
We used temporal dynamics (13) with $\boldsymbol{R}=\boldsymbol{I}-\boldsymbol{W}\boldsymbol{\Lambda}(\boldsymbol{u})$ and energy function (14) with $\boldsymbol{S}=\boldsymbol{I}$ to generate results given in Figure 1. One can also prove results similar to Proposition 5 for energy function (14) but since the extension is straightforward we do not pursue this further.
@@ 5 Conclusions
The Lyapunov function, as it is currently considered, is not of huge help. We can suggest several appropriate use cases: (i) model parameters may be adjusted on the training stage to make the Lyapunov function unstable for particular states – something that can not be easily done by integrating the dynamical system alone. This may help to learn from negative and adversarial examples [32], [6]; (ii) Lyapunov function may help in theoretical understanding of memory capacity, e.g., in the present article we have already shown that a whole flat direction corresponds to the basin of attraction, and can not support more than a single memory; (iii) Lapynov function may be directly used to compute and manipulate basins of attraction, potentially speeding up learning and making memory more robust to adversarial attacks.

### v2

@@ 4 Associative memories without dead neurons
To avoid these difficulties we will slightly modify the dynamics of associative memory (1) to minimally alter steady states and get a better energy function. In a simplified terms, our idea is to modify right-hand side of (1) such that it becomes conservative vector field. If this is the case, flat energy directions disappear. For example, the simplest model of this kind is $\dot{\boldsymbol{u}}(t)=\boldsymbol{W}^{\top}\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}(t)+\boldsymbol{b})-\boldsymbol{u}(t)=-\frac{\partial E(\boldsymbol{u})}{\partial\boldsymbol{u}}$ with energy $E(\boldsymbol{u})=\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{u}-L(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b})$. A more general dynamical system along the same lines reads:
@@
If one assumes that it is possible to select $\boldsymbol{R}(\boldsymbol{u})$ that has empty nullspace for all $\boldsymbol{u}$, i.e., that information is not lost after multiplication by $\boldsymbol{R}(\boldsymbol{u})$, the condition that we will verify later, steady states of the newly introduced system (5) follows from the old one with affine transformation $\boldsymbol{y}=\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}$. This means all architectures possible with the old model (1) are also possible with the new one (5). More precisely, the structure of steady state is the same for memory (1) and memory (5), but the structure of basins of attraction is different owing to the matrix $\boldsymbol{R}(\boldsymbol{u})$ absent in model (1).
@@
$\begin{split}&E(\boldsymbol{u};\alpha,\beta,\gamma,\boldsymbol{S})=\alpha E_{1}(\boldsymbol{u})+\beta E_{2}(\boldsymbol{u})+\gamma E_{3}(\boldsymbol{u};\boldsymbol{S}),\alpha,\beta,\gamma\in\mathbb{R},\\ &E_{1}(\boldsymbol{u})=\boldsymbol{u}^{\top}\boldsymbol{W}\boldsymbol{g}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)-L\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)-\frac{1}{2}\boldsymbol{g}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)^{\top}\boldsymbol{W}\boldsymbol{g}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right),\\ &E_{2}(\boldsymbol{u})=\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{W}\boldsymbol{u}-L\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right),\,E_{3}\left(\boldsymbol{u};\boldsymbol{S}\right)=\frac{1}{2}\left(\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b})\right)^{\top}\boldsymbol{S}\left(\boldsymbol{u}-\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b})\right)\end{split}$ (6)
@@
    $E(\boldsymbol{u};\alpha,\beta,\gamma,\boldsymbol{S})$ is a Lyapunov function for dynamical system (5) if one can find $\boldsymbol{Q}\geq 0$, $\boldsymbol{R}(\boldsymbol{u})$, $\boldsymbol{S}=\boldsymbol{S}^{\top}$ such that for all $\boldsymbol{u}$
    $\boldsymbol{R}^{\top}\boldsymbol{F}{(\boldsymbol{u})}+\boldsymbol{F}{(\boldsymbol{u})}^{\top}\boldsymbol{R}\geq\boldsymbol{Q},\boldsymbol{F}(\boldsymbol{u})=\boldsymbol{W}\boldsymbol{\Lambda}\left(\boldsymbol{W}\boldsymbol{u}+\boldsymbol{b}\right)\left(\alpha\boldsymbol{W}-\gamma\boldsymbol{S}\right)+\beta\boldsymbol{W}+\gamma\boldsymbol{S}.$ (7)
@@
Proof: Appendix E.
@@
In this case $E=E_{2}$ condition (7) becomes $\boldsymbol{R}^{\top}\boldsymbol{W}+\boldsymbol{W}\boldsymbol{R}\geq 0$. There are many strategies to select $\boldsymbol{R}$ and $\boldsymbol{W}$:
@@
Here we build a concrete example of hierarchical associative memory [15] with smooth Leaky ReLU [1] to avoid dealing with ODE having a non-smooth right-hand side. The activation function reads $\boldsymbol{g}(\boldsymbol{u})=\boldsymbol{u}\left(5+3{\sf erf}\left(3\boldsymbol{u}\big{/}8\right)\right)\big{/}8$ and its derivative is $\boldsymbol{g}^{{}^{\prime}}(\boldsymbol{u})=\boldsymbol{g}(\boldsymbol{u})\big{/}\boldsymbol{u}+9\boldsymbol{u}\exp\left(-9\boldsymbol{u}^{2}\big{/}64\right)\big{/}\left(32\sqrt{\pi}\right)$, where all operations are pointwise and ${\sf erf}$ is error function. Weights and state vector are selected as in Example 1 but without requirement $\boldsymbol{W}_{i,i+1}=\boldsymbol{W}_{i+1,i}^{\top}$. After that we select $\boldsymbol{R}=\boldsymbol{I}-\boldsymbol{W}^{\top}\boldsymbol{\Lambda}$ and obtain the following dynamical system
@@
For this system Lyapunov function $E_{2}=\frac{1}{2}\left(\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}(t)+\boldsymbol{b})-\boldsymbol{u}(t)\right)^{\top}\left(\boldsymbol{g}(\boldsymbol{W}\boldsymbol{u}(t)+\boldsymbol{b})-\boldsymbol{u}(t)\right)$ is non-increasing on trajectories. Matrix $\left(\boldsymbol{I}-\boldsymbol{W}^{\top}\odot\boldsymbol{g}^{{}^{\prime}}(\boldsymbol{W}\boldsymbol{u}(t)+\boldsymbol{b})\right)$ can in principle have non-trivial nullspace. To avoid this one observes that $\left\|\boldsymbol{W}\right\|_{2}\leq\max_{i}\left\|\boldsymbol{W}_{i,i+1}\right\|_{2}+\max_{i}\left\|\boldsymbol{W}_{i,i-1}\right\|_{2}$ (the bound is tight) and take $\boldsymbol{W}_{i,i+1}=\widetilde{\boldsymbol{W}}_{i,i\pm 1}\big{/}\left(2\left\|\widetilde{\boldsymbol{W}}_{i,i\pm 1}\right\|_{2}\right)$, for some $\widetilde{\boldsymbol{W}}$ with the same block structure. Note, that in this construction $\boldsymbol{\Lambda}$ is not positive definite, and $\boldsymbol{W}$ is not symmetric.
All theoretical statements that we derived are valid only for modified dynamical system 5. However, it is easy to see that energy function $E_{3}$ from Proposition 5 can be adopted for the original temporal dynamics (1). The resulting energy function and dynamical system reads
@@
$E_{3}({\boldsymbol{y}})=\frac{1}{2}\left(\boldsymbol{W}\boldsymbol{g}(\boldsymbol{y}(t))-\boldsymbol{y}(t)+\boldsymbol{b}\right)^{\top}\boldsymbol{S}\left(\boldsymbol{W}\boldsymbol{g}(\boldsymbol{y}(t))-\boldsymbol{y}(t)+\boldsymbol{b}\right).$ (10)
We used temporal dynamics (9) with $\boldsymbol{R}=\boldsymbol{I}-\boldsymbol{W}\boldsymbol{\Lambda}({\boldsymbol{y}})$ and energy function (10) with $\boldsymbol{S}=\boldsymbol{I}$ to generate results given in Figure 1. One can also prove results similar to Proposition 5 for energy function (10) but since the extension is straightforward we do not pursue this further.
@@ 5 Conclusions
The Lyapunov function, as it is currently considered, is not of huge help. We can suggest several appropriate use cases: (i) model parameters may be adjusted on the training stage to make the Lyapunov function unstable for particular states – something that can not be easily done by integrating the dynamical system alone. This may help to learn from negative and adversarial examples [32], [6]; (ii) Lyapunov function may help in theoretical understanding of memory capacity, e.g., in the present article we have already shown that a whole flat direction corresponds to the basin of attraction, and can not support more than a single memory; (iii) Lyapunov function may be directly used to compute and manipulate basins of attraction, potentially speeding up learning and making memory more robust to adversarial attacks.